A man has a 50-foot roll of fencing to make a rectangular shaped kennel. If he wants the kennel to be 6 feet longer than it is wide, the dimensions must be:
(A) 3.5 ft x 9.5 ft (B) 11 ft x 17 ft (C) 9.5 ft x 15.5 ft
(D) 5 ft x 11 ft (E) none of these

Answer by richwmiller(17219)

2L+2W=p
L=W+6
2(W+6)+2W=50
2W+12+2W=50
4W+12=50
4W=38
2W=19
W=9.5
L=W+6
L=15.5
(C) 9.5 by 15.5